OpenStudy (anonymous):

Simplify the following expression: 8^2 /the square root of 16 + 5^2 - 8 forty-nine twenty-one thirty-three forty-three

OpenStudy (anonymous):

anything with a ^2 after it is supposed to be squared

myininaya (myininaya):

Is that: \[\frac{8^2}{\sqrt{16+5^2-8}}\]

OpenStudy (anonymous):

\[8\div \sqrt{16}+5^{2}-8\]

myininaya (myininaya):

\[\frac{8}{\sqrt{16}}+5^2-8 \]

OpenStudy (anonymous):

so its \[1+5^{2}-8\]?

myininaya (myininaya):

\[\sqrt{16}=4\]

OpenStudy (anonymous):

or actually its \[2+5^{2}-8?\]

myininaya (myininaya):

Yes

OpenStudy (anonymous):

then you simplify the 5 squared to \[2+25-8\]

myininaya (myininaya):

yes!

OpenStudy (anonymous):

so 2+25 = 27 27-8 = 19?

myininaya (myininaya):

yes!

OpenStudy (anonymous):

then what because 19 is not an answer and im lost after that.......

myininaya (myininaya):

\[\frac{8}{\sqrt{16}}+5^2-8=\frac{8}{4}+5^2-8=2+25-8=27-8=19\] Yep it is the answer! Be sure that the problem I have looks exactly like the problem you have.

OpenStudy (anonymous):

o i see thank you alot!